# Tag&Tab: Pretraining Data Detection in Large Language Models Using Keyword-Based Membership Inference Attack

### v1

@@ Abstract
Large language models (LLMs) have become essential digital task assistance tools. Their training relies heavily on the collection of vast amounts of data, which may include copyright-protected or sensitive information. Recent studies on the detection of pretraining data in LLMs have primarily focused on sentence-level or paragraph-level membership inference attacks (MIAs), usually involving probability analysis of the target model prediction tokens. However, the proposed methods often demonstrate poor performance, specifically in terms of accuracy, failing to account for the semantic importance of textual content and word significance. To address these shortcomings, we propose Tag&Tab, a novel approach for detecting data that has been used as part of the LLM pretraining. Our method leverages advanced natural language processing (NLP) techniques to tag keywords in the input text - a process we term Tagging. Then, the LLM is used to obtain the probabilities of these keywords and calculate their average log-likelihood to determine input text membership, a process we refer to as Tabbing. Our experiments on three benchmark datasets (BookMIA, MIMIR, and the Pile) and several open-source LLMs of varying sizes demonstrate an average increase in the AUC scores ranging from 4.1% to 12.1% over state-of-the-art methods. Tag&Tab not only sets a new standard for data leakage detection in LLMs, but its outstanding performance is a testament to the importance of words in MIAs on LLMs.
@@ 1 Introduction
The rapid advancement of generative artificial intelligence (GenAI) in recent years has significantly shifted the tech industry’s focus toward the development of powerful tools such as large language models (LLMs). Used for applications ranging from conversational chats to content generation [1, 2], LLMs have become an essential tool in the digital landscape. Notable examples of foundational LLMs and their applications include OpenAI’s GPT-4 [3], which is used in applications such as ChatGPT [4], Google’s Gemini [5], which is integrated in various Google services for enhanced user interactions and AI-driven features [6], and Meta’s LLaMA [7], which is used in both research and commercial applications for diverse language understanding and content generation tasks.
@@
The rapid adoption and widespread use of these models in fields ranging from customer service to scientific research highlight the intense competition among tech companies and research institutions to develop the most advanced LLMs. While this competition drives performance improvements, it is heavily dependent on the collection of large amounts of data [8]. To achieve these improvements, LLMs are primarily trained on open-source datasets obtained from various sources using methods such as synthetic data generation and web scraping [9, 10]. The type of data collected may include books, code, academic papers, and medical records [11, 12, 13, 7]. The methods employed for data collection raise significant privacy and ethical concerns [14, 15], primarily regarding the inclusion of personally identifiable information (PII) [16] and copyright-protected content, whose inclusion is illegal, violating privacy and intellectual property rights [17, 18, 11].
Awareness regarding these privacy and ethical issues has increased as a result of legal conflicts and the lack of transparency regarding the data collection process [13, 19, 20]. The lawsuit between The New York Times and OpenAI [21], as well as other lawsuits [22, 23], highlight the critical need for mechanisms aimed at detecting such violations of privacy and intellectual property, and more specifically, identifying the data used to train LLMs [24]. Such detection can be challenging, and several techniques have been proposed for detecting LLMs’ pretraining data. Methods such as sentence-level membership inference attacks (MIAs) and paragraph-level MIAs are designed to determine whether a specific text was part of a model’s training dataset [25, 26].
To determine whether a specific piece of data was included in the training set, MIAs exploit the differences in a model’s behavior when it processes seen and unseen data. The underlying assumption is that different patterns in the LLM will be seen (e.g., higher prediction confidence or greater loss reduction) when a model encounters data it has seen during training compared to new data. Despite MIAs’ demonstrated effective performance, they have several flaws [27]. First, their performance in terms of common metrics, such as accuracy and the true positive rate (TPR) at a fixed low false positive rate (FPR), tends to worsen as the training set size increases, often rendering it close to random [28]. This is due to the tradeoff between generalization and memorization in LLMs: as the model is trained on more data, it will generalize better, while increasing the number of model parameters increases the model’s tendency to memorize the pretraining data [29, 30]. The second flaw of existing MIA methods is their reliance solely on calculating token probabilities for sections (or all) of a text. This approach fails to consider the contextual and semantic aspects of the text, which contains crucial information about text membership [31, 32]. Third, most MIAs exhibit inconsistent performance, excelling on certain models and datasets while failing to generalize across different models and data distributions [27, 24]. Lastly, MIAs are often evaluated on data that is not independently and identically distributed (IID), which can lead to the detection of distribution shifts rather than genuine membership inference, thereby undermining the attacks’ reliability [33].
@@
To address the limitations of existing methods, we introduce Tag&Tab, a novel approach based on common natural language processing (NLP) methods which is designed to efficiently and effectively detect LLMs’ pretraining data. Specifically, our method aims to determine whether an LLM was trained on a given text sample, given black-box access to the target LLM (i.e., can only query the model). Similar to other previous studies, we also assume partial knowledge of the pretraining data, i.e., word distribution [25, 34].
Inspired by the research of Lukas et al. [16], our Tag&Tab method leverages common NLP techniques for keyword selection and their subsequent prediction by the inspected LLM. Their work demonstrates the importance of focusing on named entity recognition (NER) words in identifying PII leakage, an insight which influenced the design of our method and specifically the priority it places on selecting high-entropy words in a given text. Our method consists of the following steps:
@@
    Tagging - Identifying the high-entropy words in the text and selecting the $K$ words with the highest entropy value, referred to as keywords.
@@
    Tabbing - Passing the entire text to the target LLM and calculating the average log-likelihood of these $K$ keywords.
@@
    Inference - Comparing the average log-likelihood to a threshold to determine the text’s membership (i.e., whether it was in the inspected model training set).
The intuition behind our method is that a higher log-likelihood for these challenging-to-predict high-entropy keywords indicates that the model previously encountered the text during its training phase [29]. Based on this intuition, we hypothesize that these rare keywords are more likely to be memorized by the model and thus serve as effective indicators of the text’s membership in the pretraining dataset [35, 28]. Furthermore, by selecting just a small number of high-entropy keywords, our method captures the most informative parts of the text, minimizing noise from other words’ probabilities.
We evaluated our method on nine LLMs of varying sizes and across three datasets containing 8 types of textual data. Our results demonstrate Tag&Tab’s superior performance over the examined state-of-the-art (SOTA) MIAs, with our method achieving an AUC score improvement of over 10% compared to the best-performing SOTA attack across multiple types of textual data.
The contributions of our paper are as follows.
  • We propose Tag&Tab, a novel approach for the detection of LLMs’ pretraining data that focuses on the contextual and semantic relevance of the words in a text, and opens the door to additional research on MIAs against LLMs.
  • To the best of our knowledge, we are the first to create a robust reference-free MIA method that demonstrates high and consistent performance across multiple textual data types and LLMs, while outperforming SOTA methods.
  • Our proposed approach proves to be both resource and time efficient, unlike reference-based attacks, which necessitate training a reference model, and other reference-free attacks that require an additional model [34].
Figure 1: Illustration of the Tag&Tab method - The process starts by inputting a text (in this example, the input is the conclusion of the well-known poem “The Road Not Taken" [36]) in the target LLM to obtain its word probability distribution (word probability). In the tag step, the keywords are selected based on the words’ entropy value (created in the preprocessing phase). In the tab step, the log-likelihood of the selected keywords is calculated. Finally, in the infer step, the average log-likelihood of the chosen keywords is compared against a threshold $\gamma$ to determine if the text was part of the target LLM’s pretraining data.
@@ 2 Related Work
Membership Inference (MI) [37] is a classification task that aims to determine whether a data sample $x$ was part of the training data $D_{\text{train}}$ of a model $f$. The attacker receives a sample $x$ and a model $f$, and uses the attack model $A$ to classify $x$ as a member $x\in D_{\text{train}}$ if $A(f(x))=1$; otherwise, $x$ is classified as a non-member $x\notin D_{\text{train}}$.
Large Language Model Membership Inference is a subdomain of membership inference that has received increasing research attention. Within this subdomain, the topic of detecting pretraining data has been the focus of numerous studies exploring various methodologies for identifying whether specific texts were included in the training dataset of an LLM. Existing MIAs against LLMs can be categorized as reference-free or reference-based.
Reference-based [37] methods infer data membership of a given input text by comparing the target model’s outputs with the output of reference models, which were usually trained on the same data distribution. An example of such a method is LiRA [26] which estimates the likelihood ratio of a target example’s loss under models trained with and without the example, using Gaussian distributions to simplify the computation.
In contrast, reference-free methods aim to infer membership by applying different calculation methods on the probabilities of the token predictions. Among these methods is the LOSS Attack [31], which uses the loss values of a model; in the context of language models, this corresponds to the perplexity of the text. Perplexity is a measurement of how well a probability model predicts a sample, and it is calculated as the exponentiation of the negative average log-likelihood per token:
@@
where $N$ is the number of tokens, and $P(t_{i}\mid t_{1},\ldots,t_{i-1})$ is the conditional probability of the $i$-th token given its preceding tokens. The attack relies on the assumption that a lower perplexity indicates that the text is more familiar to the model, suggesting it may be part of the training set. The Zlib attack, as presented in the paper of Carlini et al.  [32], infers membership based on calculating the ratio between the log-likelihood of the text’s and its Zlib compression length. Newer attacks, such as the Neighbor attack[34], use a different language model than the target model in order to change selected words from a given text to create new sentences, noted as neighbors’ sentences, and then calculate the ratio between the original text’s perplexity and the perplexity of its neighbors’ sentences. Although the Neighbor attack showed some success, its computational cost is very high compared to other known methods. Attacks that outperform the Neighbor attack, with low computational requirements, are the Min-K[38], and Min-K%++ [39] methods which focus on the least confident predictions from the model’s output. Min-Kcalculates the average of the lowest kprobabilities from the model’s output, and Min-K%++ extends this by normalizing token log probabilities using the mean and variance. Lastly, two recently published attacks are RECALL [40], which measures the relative change in log-likelihood when conditioning the target text on non-member prefixes, and DC-PDD [41], which calibrates token probabilities using divergence from a reference corpus, effectively mitigating the impact of high-frequency tokens. Although each of the above attacks has been shown to be successful on some datasets and models, their performance was inconsistent across the examined datasets and models in each study. However, recent research performed by Maini et al. [24] showed that aggregating the results of multiple MIAs improves the accuracy of dataset membership inference. Although this is promising, based on their findings, we conclude that the success of this aggregated approach and its use in real-world scenarios relies on improved MIAs, and specifically, attacks capable of obtaining AUC results greater than 50%.
Common Models for evaluating MIAs on LLMs include open-source models with known pretraining data. Among these models is LLaMa 1 [7], created by Meta, which used the Pile dataset training data according to Meta’s publications. Recent research has also focused on the open-source series of models known as Pythia [42]. This suite of eight LLMs, ranging in size from 70M to 12B parameters, was trained on data including the Pile dataset, with all models processing the public data in the same order during their training phase.
@@ 3 Method
We introduce Tag&Tab, a novel method for identifying data used to pre-train LLMs, which is designed to be both resource- and time-efficient. Tag&Tab leverages common NLP techniques to tag keywords in the pretraining data and predict their occurrence using the target LLM.
Our method is based on the ability to strategically select words from the input that should be challenging for the LLM to predict. Successful prediction may imply that the model previously learned the input’s content during the pretraining process. In Tag&Tab, words are selected according to their entropy values. A high entropy value indicates that a word is less common in the input text compared to other words. Knowing that LLMs tend to memorize certain parts of the pretraining data that are rare or unique, the high-entropy words are more likely to be memorized [29], particularly in the context of their preceding words. Therefore, the LLM’s predictions for these words, given the preceding context, are likely to yield higher probabilities if the model has been trained on them, compared to other high-entropy words appearing in a different context that the model was not exposed to during the training.
Tag&Tab selects $K$ words with the highest entropy value in a sentence, which are referred to as Keywords. By selecting a small number of informative keywords, we aim to capture the semantic importance of the pretraining content. This approach hinges on a hypothesis that only $K$ selected keywords are needed to accurately predict the membership of the entire input text in the pretraining data, minimizing noise from other words in the input text [38].
Tag&Tab operates under black-box constraints regarding the target model, meaning that we can produce output from a given input text and observe the token probability of the model’s predictions, but we do not have access to the model’s weights [43, 25, 34]. Additionally, we assume that prior knowledge about part of the pretraining data or its distribution is known [7, 42, 44].
@@
    Preprocessing - First, the word entropy map is constructed using the Python package wordfreq [45], which provides frequency estimates for words in a specified language. The entropy for each word is calculated using the formula:
    $E(w_{i})=p(w_{i})\cdot\log_{2}p(w_{i})$
    In the preprocessing stage, the text is also split into individual sentences, using segmentation tools (e.g., the NLTK package [46]). To avoid selecting less informative keywords due to insufficient sentence length, sentences with fewer than a specified number of words are filtered out.
@@
    Tagging - From each sentence $S$ in the text file $T$, $K$ keywords are selected, targeting named entities and words with high entropy values. This selection is based on the word entropy map for high-entropy words, and named entities are identified during this process. Tools like spaCy [47] can be used to identify named entities.
@@
    Tabbing - This stage mimics the auto-completion feature found in interfaces like a command line, where it predicts and fills in the rest of the command based on the context of the preceding input. Using the target model $M$, we infer the log-likelihood of the entire text, and then the method calculates only the previously identified keywords. For each sentence $S\in T$ consisting of $n$ words $w_{1},w_{2},\ldots,w_{n}$, where each word $w_{i}$ is decomposed into tokens, denoted as $w_{i}=t_{i_{1}},t_{i_{2}},\ldots,t_{i_{m}}$,token $t_{i_{j}}$, given its preceding tokens, is calculated as $\log p_{M}(t_{i_{j}}|t_{i_{1}},\ldots,t_{i_{j-1}})$. We define the log-likelihood of a word $w_{i}$ using the log-likelihood of its first token $t_{i_{1}}$ given its preceding tokens, expressed as $\log p_{M}(t_{i_{1}}|t_{1_{1}},t_{1_{2}},\ldots,t_{i-1_{j}})$. As a result, we obtain:
@@
[…]
@@
    Inferring - In this stage, the method calculates the average of the keywords’ probability of all the sentences in the text $T$ and compares this value to a predetermined threshold $\gamma$, to determine the text’s membership.

### v2

@@ Abstract
Large language models (LLMs) have become essential tools for digital task assistance. Their training relies heavily on the collection of vast amounts of data, which may include copyright-protected or sensitive information. Recent studies on detecting pretraining data in LLMs have primarily focused on sentence- or paragraph-level membership inference attacks (MIAs), usually involving probability analysis of the target model’s predicted tokens. However, these methods often exhibit poor accuracy, failing to account for the semantic importance of textual content and word significance. To address these shortcomings, we propose Tag&Tab, a novel approach for detecting data used in LLM pretraining. Our method leverages established natural language processing (NLP) techniques to tag keywords in the input text, a process we term Tagging. Then, the LLM is used to obtain probabilities for these keywords and calculate their average log-likelihood to determine input text membership, a process we refer to as Tabbing. Our experiments on four benchmark datasets (BookMIA, MIMIR, PatentMIA, and the Pile) and several open-source LLMs of varying sizes demonstrate an average increase in AUC scores ranging from 5.3% to 17.6% over state-of-the-art methods. Tag&Tab not only sets a new standard for data leakage detection in LLMs, but its outstanding performance is a testament to the importance of words in MIAs on LLMs.
@@ 1 Introduction
The rapid advancement of generative artificial intelligence (GenAI) in recent years has significantly shifted the tech industry’s focus toward the development of powerful tools such as large language models (LLMs).
LLMs are now widely used for tasks such as conversational AI, content generation, and scientific research [1, 2, 3, 4]. Their adoption reflects a broader shift in AI towards large-scale language understanding.
The widespread use of LLMs has intensified competition to improve model performance, which relies on the collection of vast amounts of data [5].
@@
To achieve these improvements, LLMs are primarily trained on open-source datasets obtained from various sources using methods such as synthetic data generation and web scraping [6, 7]. The type of data collected includes books, code, academic papers, and medical records [8, 9, 10, 4]. The methods employed for data collection raise significant privacy and ethical concerns [11, 12], primarily regarding the inclusion of personally identifiable information (PII) [13] and copyright-protected content [14, 15, 8].
High-profile lawsuits, such as The New York Times Company vs. OpenAI [16], highlight the need for tools that can detect unauthorized use of data in LLM training [17].
Membership inference attacks (MIAs) aim to identify whether a given text was part of a model’s training data by exploiting behavioral differences in how LLMs process seen versus unseen data (e.g., higher prediction confidence or lower loss) [18, 19]. Existing MIAs face several key limitations. First, most methods rely solely on token-level probabilities, neglecting the semantic importance of words within the broader context [20, 21]. Second, their performance varies widely across different models and datasets, often lacking consistent generalization [22, 17]. Lastly, MIAs are often evaluated on data that is not independently and identically distributed (IID), which can lead to the detection of distribution shifts rather than genuine membership inference, thereby undermining the attacks’ reliability [23].
@@
To address the limitations of existing methods, we introduce Tag&Tab, a novel approach based on common natural language processing (NLP) methods that is designed to efficiently and effectively detect LLMs’ pretraining data. Specifically, our method aims to determine whether an LLM was trained on a given text sample, given black-box access to the target LLM (i.e., can only query the model).
Building on the work of Lukas et al. [13], who highlighted the role of named entities in PII leakage detection, Tag&Tab prioritizes informative keywords using entropy-based selection.
Tag&Tab is designed to address the three key limitations of prior MIA methods. First, rather than relying solely on token-level probabilities, our method introduces semantic awareness by prioritizing meaningful content through keyword selection. Second, our results demonstrate strong generalization across models and datasets, addressing the issue of model inconsistency. Finally, while no MIA is entirely immune to distribution shifts, our focus on rare and informative keywords, rather than shallow statistical artifacts, provides better resilience to distributional variations.
Our method consists of the following steps:
@@
    Tagging - Combining (i) Named Entity Recognition (NER) spans, and (ii) the $K$ words in the sentence with the lowest entropy contribution value. The final set is not constrained to exactly $K$ words.
@@
    Tabbing - Passing the entire text to the target LLM and calculating the average log-likelihood of the selected keywords.
@@
    Inference - Comparing the average log-likelihood to a threshold to determine the text’s membership (i.e., whether it was in the inspected model’s training set).
Words with a low entropy contribution are usually either very common or very rare. After constructing the entropy map, we observed that rare content words stand out more in the low-entropy pool, making them more likely to be chosen as keywords.
Our method is based on the intuition that a higher log-likelihood for low-entropy words, together with NER spans, suggests that the model encountered the text during training [24] Based on this intuition, we hypothesize that these rare low-entropy words and rare entities are more likely to be memorized by the model and thus serve as effective indicators of the text’s membership in the pretraining dataset [25, 26]. By selecting a small number of low-entropy keywords and augmenting them with NER spans, our method captures the most informative text elements while minimizing noise from other word probabilities.
We evaluated our method on ten LLMs of varying sizes and four datasets containing nine types of textual data. Our results show that Tag&Tab outperforms state-of-the-art (SOTA) MIAs, achieving an average increase in AUC scores ranging from 5.3% to 17.6% compared to the best-performing SOTA method on multiple textual data types.
The contributions of our paper are as follows:
  • We propose Tag&Tab, a novel approach for the detection of LLMs’ pretraining data that focuses on the contextual and semantic relevance of the words in a text and opens the door to additional research on MIAs against LLMs.
  • To the best of our knowledge, this is the first robust reference-free MIA method to achieve consistently high performance across multiple textual data types and LLMs, outperforming SOTA methods.
  • Our approach is both resource- and time-efficient. Unlike reference-based attacks that require training a separate model or reference-free methods that depend on auxiliary models (e.g., the Neighbor attack [27]), Tag&Tab operates without any additional model training or inference. This minimizes computational overhead and simplifies deployment in real-world scenarios.
Figure 1: Illustration of the Tag&Tab method - The process starts by inputting a text (in this example, the input is the conclusion of the well-known poem “The Road Not Taken" [28]) in the target LLM to obtain its word probability distribution (word probability). In the tag step, the keywords are selected based on the lowest-entropy words (computed in the preprocessing phase) and NER spans. In the tab step, the log-likelihood of the selected keywords is calculated. Finally, in the infer step, the average log-likelihood of the chosen keywords is compared against a threshold $\gamma$ to determine if the text was part of the target LLM’s pretraining data.
@@ 2 Related Work
Membership inference (MI) [29] is a classification task that determines whether a data sample $x$ was part of a model’s training $D_{\text{train}}$ of a model $f$. An attacker receives a sample $x$ and a model $f$, and applies an attack model $A$ to classify $x$ as a member $x\in D_{\text{train}}$ if $A(f(x))=1$; otherwise, $x$ is classified as a non-member $x\notin D_{\text{train}}$.
Large language model membership inference is a subdomain of membership inference that has gained increasing research attention. Within this subdomain, detecting pretraining data has been the focus of numerous studies exploring different methodologies for determining whether specific texts were included in an LLM’s training dataset. Existing MIAs for LLMs fall into two categories: reference-based and reference-free.
Reference-based [29] methods compare a target model’s outputs to those of reference models, which are typically trained on the same data distribution. One such method is LiRA [19], which estimates the likelihood ratio of a target example’s loss under models trained with and without the example, using Gaussian distributions to simplify the computation.
In contrast, reference-free methods aim to determine membership by applying different probability-based calculations on token predictions. One such method, the LOSS Attack [20], uses model loss values, which in language models correspond to text perplexity. Perplexity measures how well a probability model predicts a sample and is calculated as the exponentiation of the negative average log-likelihood per token:
@@
where $N$ is the number of tokens, and $P(t_{i}\mid t_{1},\ldots,t_{i-1})$ is the conditional probability of the $i$-th token given its preceding tokens. The attack assumes that lower perplexity indicates a text is more familiar to the model, suggesting it was part of the training set. The Zlib attack [21], infers membership by calculating the ratio of a text’s log-likelihood to its Zlib compression length. Newer attacks, such as the Neighbor attack [27], modify selected words in a given text using a different language model to generate ’neighbor’ sentences, then compare the original text’s perplexity to that of its neighbors. Although the Neighbor attack showed some success, its computational cost is very high compared to other known methods. More computationally efficient attacks that outperform the Neighbor attack include Min-K[30], and Min-K%++ [31], which focus on the least confident model predictions. Min-K% calculates the average of the lowest k% probabilities from the model’s output, and Min-K%++ extends this by normalizing token log probabilities using the mean and variance. Lastly, two recently published attacks are ReCaLL [32], which measures the relative change in log-likelihood when conditioning the target text on non-member prefixes, and DC-PDD [33], which calibrates token probabilities using divergence from a reference corpus, effectively mitigating the impact of high-frequency tokens. While each of these attacks has demonstrated success on some datasets and models, their performance remains inconsistent across different studies. However, recent research performed by Maini et al. [17] showed that aggregating the results of multiple MIAs improves the accuracy of dataset membership inference. While promising, these findings suggest that the success of this aggregated approach in real-world scenarios depends on improved MIAs, meaning attacks achieving AUC scores above 0.5.
Common models for evaluating MIAs on LLMs include open-source models with known pretraining data. One such model is LLaMA 1 [4], created by Meta, which was trained on a mixture of publicly available datasets, including certain subsets of the Pile such as Books3 and Gutenberg, according to Meta’s original paper.
The open-source Pythia model suite [34], which includes eight LLMs ranging from 70M to 12B parameters. These models were trained on data including the Pile dataset, with all models processing the public data in the same order during training.
@@ 3 Method
We introduce Tag&Tab, a novel resource- and time-efficient method for identifying data used to pretrain LLMs. Tag&Tab applies common NLP techniques to tag keywords in the pretraining data and predict their membership using the target LLM.
Our method strategically selects words from the input that should be challenging for the LLM to predict. Successful prediction may indicate that the model previously learned the content during pretraining. In Tag&Tab, words are selected according to their entropy contribution values. Words with a low entropy contribution value tend to be either very common (e.g., ’or,’ ’in’) or very rare. After constructing an entropy map, we found that the low-entropy pool tends to be dominated by rare content words, which are therefore more likely to be selected as keywords.
Tag&Tab relies on the hypothesis that while LLMs tend to memorize repeated data, they also memorize rare or unique training data. This aligns with the findings of Carlini et al. [24], who stated that: "memorization does still happen, even with just a few duplicates—thus, deduplication will not perfectly prevent leakage." This insight was further validated by Lukas et al. [13], who found that PIIs tokens, which are distinctive rare tokens, are effective in revealing memorized content. Thus, an LLM is likely to assign a higher probability to these low-entropy words when they appear in familiar contexts it has seen before, compared to the same low-entropy words in novel or unfamiliar contexts. For example, in the sentence "Astronomers detected light from the edge of the universe emitted by a quasar," the word "quasar" is a low-entropy word. If this sentence appeared in the training data, the model will assign a higher probability to "quasar" in this context than if it had not seen the sentence before.
Tag&Tab selects $K$ words with the lowest entropy values in a sentence, which are referred to as keywords. By selecting a small number of informative keywords, it aims to capture the semantic importance of the pretraining content. This approach hinges on a hypothesis that only $K$ selected keywords are needed to accurately predict the membership of the entire input text in the pretraining data, minimizing noise from other words in the input text [30].
Tag&Tab operates under black-box constraints, meaning that we can observe token probabilities from a given input but lack access to the model’s weights, which is standard practice in MIAs [35, 18, 27, 31].
@@
    Preprocessing - First, the word entropy map is constructed using the Python package wordfreq [36], which provides frequency estimates for words in a specified language. The entropy contribution value for each word is calculated using the formula:
    $H(w_{i})=-p(w_{i})\cdot\log_{2}p(w_{i})$
    This value is high for mid-frequency words and low for both very common and very rare words. In this stage, the text is also split into individual sentences, using segmentation tools (e.g., the NLTK package [37]). To avoid selecting less informative keywords due to insufficient sentence length, sentences with fewer than a specified number of words are filtered out.
@@
    Tagging - From each sentence $S$ in the text file $T$, the method first takes the $K$ lowest-entropy words according to the entropy map. We then identify named-entity spans (e.g., people, organizations, locations) using spaCy [38]. The final keyword set is the union of the low-entropy words and the NER spans. Because this is a union, the set may be larger than $K$, but in practice it is often exactly $K$, because NER spans often overlap with low‑entropy tokens. For clarity, we continue to denote the size of this final set as $K$, referring to the number of keywords after the union.
@@
    Tabbing - This stage mimics the auto-completion feature found in interfaces like a command line, where it predicts and fills in the rest of the command based on the context of the preceding input. Using the target model $M$, the method computes the log-likelihood of the entire text, then focuses on the log-likelihood of the previously identified keywords. For each sentence $S\in T$ consisting of $n$ words $w_{1},w_{2},\ldots,w_{n}$, where each word $w_{i}$ is decomposed into tokens, denoted as $w_{i}=t_{i_{1}},t_{i_{2}},\ldots,t_{i_{m}}$, token $t_{i_{j}}$, given its preceding tokens, is calculated as $\log p_{M}(t_{i_{j}}|t_{i_{1}},\ldots,t_{i_{j-1}})$. We define the log-likelihood of a word $w_{i}$ using the log-likelihood of its first token $t_{i_{1}}$ given its preceding tokens, expressed as $\log p_{M}(t_{i_{1}}|t_{1_{1}},t_{1_{2}},\ldots,t_{i-1_{j}})$. As a result, we obtain:
@@
    $\text{Keywords' Prob}(S)=\frac{1}{K}\sum_{w_{i}\in\text{Keywords}(S)}\log p_{M}(w_{i}|\cdot)$
@@
    Inferring - In this stage, the method calculates the average probability of the keywords across all sentences in text $T$ and compares it to a predetermined threshold $\gamma$ to determine membership.